Fermi National Laboratory


Accelerator Update

Friday October 15

Tevatron

    Bo Low Beta Separators HLS Installation - 25% complete

    D0 Low Beta Separators HLS Installation - 25% complete

    A0 Vacuum Upgrade - 90% complete

    IPM installation - 50% complete. The Crystal Collimator has been installed

    D4 Spool Replacement - 80%

    D17 Separator Installation - 90% complete

    Shimming 412 Dipoles - 75% complete

    Unrolling Magnet - 80% complete

      D0 Separator unroll - 65% complete
      B0 Separator unroll - 0% complete

Linac

    The Linac began sending beam to its dump at 1:10 PM on Wednesday, October 13, 2004.

Booster

    The Booster should be up and running beam by Wednesday, October 20, 2004.

    RF work completion waiting on tunnel cleaning.

    Long 13 work complete except for flow switches (Long 13 is a sector of the Booster ring.)

Antiproton Source

    Debuncher Injection Area Beampipe and Quadrupole (Quad) replacement - Checkout in progress. Work on the last four D40 quads are in progress. Septum magnet pulse tested at full current. The beam pipe is almost ready. The kicker girder and cable pulse in progress

      Debuncher Motorized Quads Installation - ~90% complete

    Accumulator Flying Wires - The wires have been installed and tested. They are waiting on vacuum pump down.

    ARF1-1 Installation - ~95% complete. Alignment completed.

    AP40 - Replaced vertical flying wire. Bake out ends today.

    AP10 SO Cord Remediation - AP10, F23, and F27 work complete. AP50 work begins no Monday.

    D/A Lines - alignment check complete, cooling system measurements next week, visual inspections after "as founds", and septum pulse testing has begun

    Beam Lines - Work in progress: AP1 magnet moves new VT108 & reconfigure vacuum windows, AP2 upstream vacuum will be opened next week.

    Other Work - Target Station, Alignment, LCW, and LLRF - all work on track

Main Injector/Recycler

    Overall shut down progress - 80% complete

    MI LCW system is up and running.

    MI-30

      Q307-309 Vacuum Modifications- 7% complete
      Q303-305 Vacuum Modifications - 100% complete

    MI-30 & 62

    MI-31

      Electron Cooling overall - 53% complete

    MI-10

      Cooling Tank Repairs - ~78% complete

MiniBooNE

    Horn Replacement - 61%, the old Horn has been moved to TSB. New horn installation scheduled for 10/20/04

NuMI

    The kicker is pulsing. The Recycler shielding is completed. The trim magnets await LVW hookup. Alignment is catching up to its schedule. Power supply test is occurring during the evening


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/19/2004   email Fermilab